### v2

# The Kuramoto model on the Sierpinski Gasket II: Twisted states

Abstract
We study the Kuramoto model (KM) of coupled phase oscillators on graphs approximating the Sierpinski gasket (SG). As the size of the graph tends to infinity, the limit points of the sequence of stable equilibria in the KM correspond to the minima of the Dirichlet energy, i.e., to harmonic maps from the SG to the circle. We provide a complete description of the stable equilibria of the continuum limit of the KM on graphs approximating the SG, under both Dirichlet and free boundary conditions. We show that there is a unique stable equilibrium in each homotopy class of continuous functions from the SG to the circle. These equilibria serve as generalizations of the classical twisted states on ring networks. Furthermore, we extend the analysis to the KM on post-critically finite fractals. The results of this work reveal the link between self-similar organization and network dynamics.
2020 Mathematical Subject Classification. 34C15, 28A80, 37B35, 37M15, 58E30, 92B20.
Keywords and phrases. Kuramoto model, $XY$ spin model, fractal, Sierpinski Gasket, $\Gamma$-convergence, twisted state
1 Introduction
In this paper, we initiate the study of the Kuramoto model (KM) of coupled phase oscillators on self-similar networks. This model has been extensively studied in nonlinear science [20] and in statistical physics, where it is known under the name of $XY$-model [7]. In particular, it has been used for understanding the role of the network topology on emergent dynamics and synchronization [5]. In a similar vein, we use the KM to explore how self-similar network organization influences network dynamics. There is empirical evidence of self-similarity in natural and technological networks [12]. Often, this is a by-product of a network’s hierarchical structure, as seen in the Internet, where smaller modules mirror the structure of the larger network [9]. While self-similarity in real-life networks is not exact from a mathematical standpoint, the analysis of idealized models on fractals can offer insights into the implications of self-similarity observed in such networks.
As a model of self-similar connectivity, we use a sequence of graphs approximating a fractal domain $K\subset\mathbb{R}^{d}$. To simplify presentation, we first focus on the case when $K$ is a Sierpinski Gasket (SG), a canonical example of a fractal set [10, 19]. In Section 7, we will extend this analysis to the KM on graphs approximating a post-critically finite (p.c.f.) fractal.
Until Section 7, $K\subset\mathbb{R}^{2}$ is the unique nonempty compact set satisfying the following fixed point equation
$K=\bigcup_{i=1}^{3}F_{i}(K),\qquad F_{i}(x)=2^{-1}(x-v_{i})+v_{i},$ (1.1)
where $v_{1},v_{2},v_{3}$ are three distinct points in the plane (see Fig. 1). Without loss of generality, we set $v_{1}=(0,0),$ $v_{2}=(1/2,\sqrt{3}/2)$, and $v_{3}=(1,0)$, so that $v_{1},v_{2},$ and $v_{3}$ are vertices of an equilateral triangle with side length $1$.
Figure 1: SG.
The sequence of graphs $\Gamma_{n}=(V_{n},E_{n})$ approximating $K$ is constructed as follows. Let $V_{n}$ stand for the set of vertices of $\Gamma_{n}$. Then $V_{0}=\{v_{1},v_{2},v_{3}\}$ and for $n\geq 1$ set
$V_{n}=\bigcup_{i=1}^{3}F_{i}(V_{n-1}).$
Further, $\Gamma_{0}$ is the complete graph on three nodes. To describe the set of edges of $\Gamma_{n}$, we need an alphabet $S\doteq\{1,2,3\}$ and the set of words consisting of $n$ symbols, $S^{n}$. Then $v_{i},v_{j}\in V_{n}$ are adjacent (denoted by either $v_{i}\sim_{n}v_{j}$, or simply $i\sim_{n}j$) if there is $w=(w_{1},w_{2},\dots,w_{n})\in S^{n}$, such that
$v_{i},v_{j}\in F_{w}(V_{0}),\quad\mbox{where}\quad F_{w}\doteq F_{w_{1}}\circ F_{w_{2}}\circ\dots\circ F_{w_{n}}.$
Geometrically, $\Gamma_{0}$ is a triangle with vertices $v_{1},v_{2},$ and $v_{3}$, and $\Gamma_{n}=\bigcup_{w\in S^{n}}F_{w}(\Gamma_{0})$. Two nodes $v_{i},v_{j}\in V_{n}$ are adjacent if both belong to the same nn–cell $F_{w}(\Gamma_{0})$ for some $w\in S^{n}$ (see Fig. 2).
Figure 2: Graphs $\Gamma_{n}$ approximating SG.
There are $\frac{3}{2}\left(3^{n}-1\right)$ vertices of $\Gamma_{n}$. We will distinguish between the boundary vertices given by $V_{0}=\{v_{1},v_{2},v_{3}\}\subset V_{n}$ and the interior vertices $V_{n}\setminus V_{0}$. Note that the interior vertices have degree $4$ while the boundary ones have degree $2$.
We pause to explain the coding of the nodes of the graphs approximating $K$. Let $T$ denote the solid triangle with vertices $v_{1},v_{2},$ and $v_{3}$. Every node from $V_{n}$ is a vertex of the corresponding triangle
$T_{w}=F_{w}(T),\;w=(w_{1},w_{2},\dots,w_{n})\in S^{n}$
and can be represented as
$\bigcap_{k=1}^{\infty}F_{w\underbrace{\scriptscriptstyle iii\dots i}_{\text{$k$ times}}}\mskip 3.0mu(T).$
For each node in $V_{n}$ we define an itinerary $w\bar{i}$, where $\bar{i}$ stands for the infinite sequence of $i$’s: $iii\dots,$ $i\in S$. Note that for $v\in V_{n}\setminus V_{0}$ there are two possible itineraries, e.g., $v_{1\bar{2}}$ and $v_{2\bar{1}}$ correspond to the same node from $V_{2}$. We order the vertices in $V_{n}$ by the corresponding itineraries in lexicographical order:
$V_{n}=\{v_{\bar{1}},v_{\bar{2}},v_{\bar{3}},v_{1\bar{1}},v_{1\bar{2}},\dots\}.$ (1.2)
Abusing notation, we will use $v_{1},v_{2},v_{3},v_{4},\dots$ to denote the elements of $V_{n}$ appearing in the same order as in (1.2).
Now we are ready to formulate the KM on $\Gamma_{n}$:
$\dot{u}(t,v_{i})=\left(\frac{5}{3}\right)^{n}\sum_{j\sim_{n}i}\sin\left(2\pi\left(u(t,v_{j})-u(t,v_{i})\right)\right),\quad v_{i}\in V_{n},$ (1.3)
where $u(t,v_{i})$ represents the state of oscillator located at $v_{i}$ at time $t$. The scaling factor $\left(\frac{5}{3}\right)^{n}$ is used so that the model has a nontrivial continuum limit as $n\to\infty$.
It is instructive to rewrite the KM on $\Gamma_{n}$ as the gradient system
$\dot{u}=-2\pi\nabla{\mathcal{J}}_{n}(u),\quad u=(u(t,v_{1}),u(t,v_{2}),\dots,u(t,v_{k_{n}})),$ (1.4)
where
${\mathcal{J}}_{n}(u)\doteq\left(\frac{5}{3}\right)^{n}\frac{1}{4\pi^{2}}\sum_{(i,j)\in E_{n}}\left(1-\cos\left(2\pi(u(t,v_{j})-u(t,v_{i}))\right)\right).$ (1.5)
We use the convention that each (undirected) edge $(i,j)\in E_{n}$ appears only once in the sum to avoid double-counting. The factor $\frac{1}{4\pi^{2}}$ is taken for convenience of future analysis.
Since (1.4) is a gradient system, its attractor coincides with the set of minima of the energy function ${\mathcal{J}}_{n}$. Due to translation invariance of (1.3) the points of minimum of ${\mathcal{J}}_{n}$ are not isolated. To eliminate this translation invariance, we fix the value at $v_{1}$ in the steady-state solutions of (1.3) to 0. This leads to the following minimization problem:
${\mathcal{J}}_{n}\;\rightarrow\;\min_{H_{n}},\quad H_{n}\doteq\{u\in L(V_{n},\mathbb{T}):\;u(v_{1})=0\},$ (1.6)
where $L(X,Y)$ denotes the space of functions with domain $X$ and co-domain $Y$.
The main result of this paper reveals the link between the structure of the attractor of (1.3) and the topology of the SG. We show that there is one-to-one correspondence between the homotopy classes of continuous functions from the SG to $\mathbb{T}$ and stable equilibria of (1.3).
Before formulating the main result, we review the homotopy classes of continuous maps on SG following [14]. To this end, let
$\mathcal{P}_{n}=\left\{\partial T_{w},\;w\in S^{n}\right\}\quad\mbox{and}\quad\mathcal{P}=\bigcup_{n=0}^{\infty}\mathcal{P}_{n}.$
For each triangular loop $\gamma\in\mathcal{P}$, we choose a reference point $O_{\gamma}\in\gamma$. For concreteness, let $O_{\gamma}$ be the leftmost vertex of $\gamma$. Choose a uniform parametrization $c_{\gamma}:\mathbb{T}\to\gamma$, which starts at $O_{\gamma}$, $c_{\gamma}(0)=O_{\gamma}$, and traces $\gamma$ in clockwise direction with constant speed.
For a given $f\in C(K,\mathbb{T})$, $f_{\gamma}\doteq f\circ c_{\gamma}$ is a continuous function from $\mathbb{T}$ to itself [^1] Throughout this paper, $c_{\gamma}$ is a uniform parametrization, i.e., $\dot{c}_{\gamma}=1.$. There is a unique continuous function $\bar{f}_{\gamma}:\mathbb{R}\to\mathbb{R}$ such that
$\hat{f}_{\gamma}(0)$ $=f(0),$ (1.7)
$\hat{f}_{\gamma}(x)$ $=f(x\mod 1)+k(x),\quad k(x)\in\mathbb{Z}.$ (1.8)
The second condition (1.8) can be written as
$\pi\circ\hat{f}_{\gamma}=f_{\gamma}\circ\pi,$ (1.9)
where $\pi:\mskip 3.0mu\mathbb{R}\ni x\mapsto x\mod 1.$
$\hat{f}_{\gamma}$ is called the lift of $f_{\gamma}$. The degree of $f_{\gamma}$ is expressed in terms of the lift of $f_{\gamma}$:
$\omega(f_{\gamma})\doteq\hat{f}_{\gamma}(1)-\hat{f}_{\gamma}(0).$
Definition 1.1.
The degree of $f\in C(K,\mathbb{T})$ is defined by
$\bar{\omega}(f)=\left(\omega_{\gamma_{0}}(f),\omega_{\gamma_{1}}(f),\omega_{\gamma_{2}}(f),\dots\right),$ (1.10)
where $\omega_{\gamma}(f)\doteq\omega(f_{\gamma})$.
Remark 1.2.
Since $f$ is uniformly continuous on $K$, the number of nonzero entries in (1.10) is finite. Unless specified otherwise, we use the topology on $K$ induced by the Euclidean metric.
Definition 1.3.
Two maps $f,g\in C(K,\mathbb{T})$ are called homotopic, denoted $f\sim g$, if there exists a continuous mapping $F:[0,1]\times K\to\mathbb{T}$ such that
$F(0,\cdot)=f\qquad\mbox{and}\quad F(1,\cdot)=g.$ (1.11)
Theorem 1.4.
(Theorem 2.4 in [14]) Let $f,g\in C(K,\mathbb{T})$. Then $f\sim g$ if and only if $\bar{\omega}(f)=\bar{\omega}(g)$.
We will need the following result on harmonic maps from SG to the circle. It follows from [17]. For a given $\omega^{\ast}\in\mathbb{Z}^{\ast}\doteq\bigcup_{l=1}^{\infty}\mathbb{Z}^{l}$ there exists a unique solution $u^{\ast}\in L(K,\mathbb{T})$ of the following boundary value problem:
$\Delta u^{\ast}$ $=0,$ (1.12)
$u^{\ast}(v_{1})$ $=0,$ (1.13)
$\partial_{\bf n}u^{\ast}(v_{i})$ $=0,i=1,2,3,$ (1.14)
$\bar{\omega}(u^{\ast})$ $=\omega^{\ast},$ (1.15)
where $\partial_{\bf n}$ stands for the normal derivative [18].
We can now formulate the main result of this work. Below, we will generalize this result for the KM on p.c.f. fractals, but for now we restrict our attention to the KM on SG.
Theorem 1.5.
Let $\omega^{\ast}\in\mathbb{Z}^{\ast}$ and $u^{\ast}\in L(K,\mathbb{T})$ be the solution of the boundary value problem (1.12)-(1.15). Then, for each $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for all $n\geq N$, the KM on $\Gamma_{n}$ has a stable steady state solution $u^{n}\in L(V_{n},\mathbb{T})$ such that
$\max_{x\in V_{n}}|u^{n}(x)-u^{*}(x)|<\epsilon.$ (1.16)
Moreover, there is an extension of $u^{n}$ to a continuous function on $K$, $\tilde{u}^{n}$ such that $\omega(\tilde{u}^{n})=\omega^{\ast}$ and
$\max_{x\in K}|\tilde{u}^{n}(x)-u^{*}(x)|<\epsilon.$ (1.17)
Figure 3 shows two examples of stable equilibria of KM on SG with non-trivial degrees. The KM equilibria in Theorem 1.5 are analogous to twisted states of the KM on nearest-neighbor graphs (cf. [22]; see also Section 4 for more details). Distinct twisted states are homotopic if and only if they share the same winding number $q$ so classifying twisted states is straightforward. On the other hand, on the fractal $K$ the homotopy of a map is determined by its degree (1.10), which can have arbitrary (finite) length. As a result, there is a rich diversity of KM equilibria of increasing topological complexity on the fractal $K$.
a) Refer to caption b) Refer to caption
Figure 3: Twisted states on SG are equilibria of the KM on SG. The degrees of the equilibria shown above: a) $\bar{\omega}(f)=(1,0,0,\dots)$, and b) $\bar{\omega}(f)=(1,1,1,1,0,0,\dots)$. These plots were obtained by numerically solving the initial value problem for the KM (1.3) on on $\Gamma_{8}$ with initial conditions taken as the corresponding harmonic map $u^{*}$ satisfying (1.12)-(1.15) and restricted to $V_{8}$.
The remainder of this paper will be mainly focused on proving Theorem 1.5. We will begin by considering a simpler variation of the problem with trivial degree (i.e., $\bar{\omega}(u^{*})=(0,0,\dots)$) and Dirichlet boundary conditions (cf. (1.14)). This problem is interesting in its own right, but also serves to build intuition and develop the necessary tools for the main result. In Section 2 we review the necessary background on harmonic functions on SG, and in Section 3 we combine $\Gamma$-convergence and a priori Hölder estimates to establish the convergence of KM equilibria to these (real-valued) harmonic functions (Theorem 3.4). In Section 4 we translate these techniques to the setting of the KM on nearest-neighbor graphs. In particular, we show how covering spaces are used to adapt the results to $\mathbb{T}$-valued harmonic maps. We will then be prepared to prove Theorem 1.5. In Section 5 we first review the necessary theory of $\mathbb{T}$-valued harmonic maps on SG, and in Section 6 we complete the remaining details for Theorem 1.5. Finally, in Section 7 we describe how these results can be applied in a more general setting to study the KM on post-critically finite fractals.
2 Harmonic functions on SG
Harmonic functions can be defined as minimizers of the Dirichlet energy (cf. [21]). On a Euclidean domain, the Dirichlet energy is given by
${\mathcal{I}}(u)=\frac{1}{2}\int_{\Omega}\left|\nabla u\right|^{2}dx,\qquad\Omega\subset\mathbb{R}^{d}.$ (2.1)
On a fractal set, partial derivatives are not defined in general. Instead, the Dirichlet energy is defined as a limit of the functionals ${\mathcal{E}}_{n}$ defined using graphs approximating a given fractal (see Figure 2). For the reader’s convenience, we outline the limiting procedure below and refer the interested reader to [10] for details. Although the approach we describe applies to a class of fractals known as p.c.f. fractals, we restrict our discussion to SG for clarity of presentation. The extension to p.c.f. fractals will be explained in Section 7.
First, we construct the sequence of graphs $(\Gamma_{m})$ approximating SG (Figure 2). On $\Gamma_{m}$ we define the discrete Laplacian
$\Delta_{m}f(v_{i})=\left(\frac{5}{3}\right)^{m}\sum_{j\sim_{m}i}\left[f(v_{j})-f(v_{i})\right],\quad f\in L(V_{n},\mathbb{R}).$ (2.2)
A function $f_{m}^{*}\in L(V_{m},\mathbb{R})$ is called $\Gamma_{m}$-harmonic if it satisfies the discrete Laplace equation at every interior node of $\Gamma_{m}$:
$\Delta_{m}f_{m}^{*}(v)=0\quad\forall v\in V_{m}\setminus V_{0}.$ (2.3)
Equivalently, harmonic functions are critical points of the associated Dirichlet forms
${\mathcal{E}}_{m}(f)\doteq\left(\frac{5}{3}\right)^{m}\sum_{(i,j)\in E_{n}}\frac{(f(v_{j})-f(v_{i}))^{2}}{2},\quad f\in L(V_{n},\mathbb{R}).$ (2.4)
The sequence of Dirichlet forms $(\mathcal{E}_{m})$ has the following properties.
  1. 1.
    A $\Gamma_{m}$-harmonic $f_{m}^{\ast}\in L(V_{m},\mathbb{R})$ minimizes $\mathcal{E}_{m}$ over all functions subject to the same boundary conditions
    $\mathcal{E}_{m}(f_{m}^{\ast})=\min\left\{\mathcal{E}(f):\;f\in L(V_{m},\mathbb{R}),\;f|_{V_{0}}=f_{m}^{\ast}|_{V_{0}}\right\}.$ (2.5)
  2. 2.
    The minimum of the energy over all extensions of $f\in L(V_{m-1},\mathbb{R})$ to $V_{m}$ is equal to $\mathcal{E}_{m-1}(f)$:
    $\min\left\{\mathcal{E}_{m}(\tilde{f}):\;\tilde{f}\in L(V_{m},\mathbb{R}),\tilde{f}|_{V_{m-1}}=f\in L(V_{m-1},\mathbb{R})\right\}=\mathcal{E}_{m-1}(f).$ (2.6)
    In particular,
    $f^{\ast}_{m+1}|_{V_{m}}=f^{\ast}_{m}.$ (2.7)
The first property follows from the Euler-Lagrange equation for ${\mathcal{E}}_{m}$. This property does not depend on the scaling coefficient $(5/3)^{m}.$ The second property, on the other hand, holds due to the choice of the scaling constant. The sequence of $(\mathcal{E}_{m})$ equips $K$ with a harmonic structure (cf. [10]).
Now let $f$ be a continuous function on $K$, $f\in C(K,\mathbb{R})$. Define $P_{m}:C(K,\mathbb{R})\to L(V_{m},\mathbb{R})$ by restricting $f$ to $V_{m}$:  $P_{m}f\doteq f|_{V_{m}}.$ We next extend the definition of ${\mathcal{E}}_{m}$ to functions in $C(K,\mathbb{R})$ (which we also denote by ${\mathcal{E}}_{m}$ by abuse of notation):
${\mathcal{E}}_{m}(f):={\mathcal{E}}_{m}\circ P_{m}(f).$
By (2.6), $({\mathcal{E}}_{m}(f))$ is a non-decreasing sequence. Thus,
${\mathcal{E}}(f)\doteq\lim_{m\to\infty}{\mathcal{E}}_{m}(f)$ (2.8)
defines a functional on $C(K,\mathbb{R})$.
The domain of the Laplacian on $K$ is defined by
$\operatorname{dom}({\mathcal{E}})=\left\{f\in C(K,\mathbb{R}):\;{\mathcal{E}}(f)<\infty\right\}.$
Definition 2.1.
A function $f\in\operatorname{dom}({\mathcal{E}})$ is called harmonic if it minimizes ${\mathcal{E}}(f)$ over all continuous functions on SG subject to boundary conditions on $V_{0}$.
The combination of (2.5) and (2.6), implies that the restriction of harmonic $f\in\operatorname{dom}({\mathcal{E}})$ to $\Gamma_{m}$, $f\left|{}_{\Gamma_{m}}\right$., is $\Gamma_{m}$-harmonic on $\Gamma_{m}$ for every $m\in\mathbb{N}$.
The second property of the energy form (cf. (2.6)) yields a recursive algorithm for computing the values of a harmonic function on $V_{\ast}=\bigcup_{m=0}^{\infty}V_{m}$, a dense subset of $K$. For $m=0$, the values on $V_{0}$ are prescribed:
$f|_{V_{0}}=\phi.$
Given $f|_{V_{m-1}},\;m\geq 1,$ the values on $V_{m}\setminus V_{m-1}$ are computed using the following $\frac{1}{5}-\frac{2}{5}$ rule, which we state for an arbitrary fixed (m1)(m-1)-cell $T_{w},\;w\in S^{m-1}$: Suppose the values of $f$ at $a,b,c$, the nodes of $T_{w}$, are known. Then the values of $f$ at $x,y,z,$ the nodes at the next level of discretization are computed as follows
$\begin{pmatrix}f(x)\\ f(y)\\ f(z)\end{pmatrix}=\begin{pmatrix}\frac{2}{5}&\frac{2}{5}&\frac{1}{5}\\ \frac{1}{5}&\frac{2}{5}&\frac{2}{5}\\ \frac{2}{5}&\frac{1}{5}&\frac{2}{5}\end{pmatrix}\begin{pmatrix}f(a)\\ f(b)\\ f(c)\end{pmatrix}$, (2.9)
see Fig. 4.
Figure 4: Harmonic extension algorithm for harmonic functions, see (2.9).
This algorithm is called harmonic extension [18]. Using (2.6), one can show that at each step of the harmonic extension:
${\mathcal{E}}_{m}(f|_{V_{m}})={\mathcal{E}}_{0}(f|_{V_{0}})=\min\left\{{\mathcal{E}}_{m}(f):\;f\in L(V_{m},\mathbb{R}),\;f|_{V_{0}}=\phi\right\}.$
Note that each of $f|_{V_{m}}$ is a $\Gamma_{m}$-harmonic function. Further, harmonic extension results in a uniformly continuous function on $V_{*}$. Thus, the values of $f$ on $G\setminus V_{*}$ can be obtained from those on $V_{*}$ by continuity.
Finally, the following Holder estimate will be used to provide compactness, which is needed for the convergence of the energies and the equilibria of the KM.
Lemma 2.2 (cf. [18, 11]).
If $f\in\operatorname{dom}({\mathcal{E}})$ then
$\frac{|f(x)-f(y)|}{|x-y|^{\beta}}\leq C\sqrt{\mathcal{E}(f)},$
where $\beta=\frac{\log(5/3)}{2\log(2)}$ and $C$ is independent of $f$.
3 $\Gamma$-convergence
$\Gamma$-convergence is our main tool for studying the asymptotic behavior of stable equilibria in the KM for large $n$. Below we recall the definition of $\Gamma$-convergence on metric spaces and refer the reader to [3] for further details.
Definition 3.1.
Let $(X,d)$ be a metric space and $F_{n},F\colon X\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$ be a sequence of functionals. Then $F_{n}$ $\Gamma$-converges to $F$, written $\Gamma$-$\lim F_{n}=F$, if the following two conditions hold.
  1. C1.
    For all $x_{n}\in X$ with $x_{n}\to x\in X$, $F(x)\leq\liminf_{n\to\infty}F_{n}(x_{n})$.
  2. C2.
    For all $x\in X$, there is some $x_{n}\to x$ so that $F(x)\geq\limsup_{n\to\infty}F_{n}(x_{n})$.
Remark 3.2.
Sequence $(x_{n})$ in C2 is called a recovery sequence. The combination of C1 and C2 implies that for every $x\in X$, there exists $x_{n}\to x$ such that
$F(x)=\lim_{n\to\infty}F_{n}(x_{n}).$
Let $X$ be the space of continuous functions on $K$, $C(K,\mathbb{R})$, equipped with the supremum metric
$d(f,g)\doteq\sup_{x\in K}|f(x)-g(x)|.$ (3.1)
First, we note that ${\mathcal{E}}_{n}$ $\Gamma$-converges to $\mathcal{E}$}. This follows from the monotone pointwise convergence of ${\mathcal{E}}_{n}\nearrow{\mathcal{E}}$ (cf. (2.8)) and continuity of ${\mathcal{E}}_{n}$ (cf. [2, Remark 1.10(ii)]). Importantly, ${\mathcal{J}}_{n}$ also $\Gamma$-converges to $\mathcal{E}$} as shown in the following theorem.
Theorem 3.3.
Let $({\mathcal{J}}_{n})$ be the Kuramoto energy functionals (cf. (1.5)). Then
${\mathcal{E}}=\Gamma\dash\lim_{n\to\infty}{\mathcal{J}}_{n}.$ (3.2)
Moreover, suppose $(u^{n})$ is a sequence of minimizers of $({\mathcal{J}}_{n})$ converging to $u^{\ast}\in X$, then $u^{\ast}$ is a minimizer of $\mathcal{E}$ and
$\lim_{n\to\infty}{\mathcal{J}}_{n}(u^{n})={\mathcal{E}}(u^{\ast}).$ (3.3)
Proof.
By the Taylor’s expansion of $\cos$,
${\mathcal{J}}_{n}(u)$ $=\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}\frac{(u(v_{j})-u(v_{i}))^{2}}{2}\left[1+g(u(v_{j})-u(v_{i}))\right]$
$={\mathcal{E}}_{n}(u)+\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}g(u(v_{j})-u(v_{i}))\frac{(u(v_{j})-u(v_{i}))^{2}}{2},$ (3.4)
where $g(x)=-\frac{2(2\pi x)^{2}}{4!}+\frac{2(2\pi x)^{4}}{6!}$-$\cdots$. Using the alternating series remainder bound, we have
$|g(x)|\leq\frac{\pi^{2}|x|^{2}}{3}.$ (3.5)
Suppose $u^{n}\to u\in\operatorname{dom}({\mathcal{E}})$. Then
$\liminf_{n}{\mathcal{J}}_{n}(u^{n})$ $\geq\liminf_{n}{\mathcal{E}}_{n}(u^{n})+\liminf_{n}\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}g(u^{n}(v_{j})-u^{n}(v_{i}))\frac{(u^{n}(v_{j})-u^{n}(v_{i}))^{2}}{2}$
$\geq\mathcal{E}(u)+\liminf_{n}\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}g(u^{n}(v_{j})-u^{n}(v_{i}))\frac{(u^{n}(v_{j})-u^{n}(v_{i}))^{2}}{2},$
where we used $\Gamma\dash\lim_{n\to\infty}{\mathcal{E}}_{n}={\mathcal{E}}$.
Using (3.4) and (3.5), we continue
$\left|\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}g(u^{n}(v_{j})-u^{n}(v_{i}))\frac{(u^{n}(v_{j})-u^{n}(v_{i}))^{2}}{2}\right|$ (3.6)
$\leq\max_{j\sim_{n}i}|g(u^{n}(v_{j})-u^{n}(v_{i}))|\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}\frac{(u^{n}(v_{j})-u^{n}(v_{i}))^{2}}{2}$
$\leq C\max_{j\sim_{n}i}|u^{n}(v_{j})-u^{n}(v_{i})|^{2}\left(\frac{5}{3}\right)^{n}\sum_{(j,i)\in E_{n}}\frac{(u^{n}(v_{j})-u^{n}(v_{i}))^{2}}{2}$
$\leq C\left[\max_{j\sim_{n}i}\frac{|u^{n}(v_{j})-u^{n}(v_{i})|^{2}}{|v_{j}-v_{i}|^{2\beta}}\right]\max_{j\sim_{n}i}|v_{j}-v_{i}|^{2\beta}{\mathcal{E}}_{n}(u^{n})$
$\leq C(\mathcal{E}(u^{n}))^{2}\max_{j\sim_{n}i}|v_{j}-v_{i}|^{2\beta},$
where we used Lemma 2.2 to obtain the bound in the last line.
Since $u^{n}\to u$, the continuity of $\mathcal{E}$ implies a uniform bound on ${\mathcal{E}}(u^{n})$. Finally, by construction of $\Gamma_{n}$,
$|v_{j}-v_{i}|=2^{-n},\quad\mbox{ when }j\sim_{n}i.$
Thus, $\max_{j\sim_{n}i}|v_{j}-v_{i}|^{2\beta}\to 0$ as $n\to\infty$, and
$\lim_{n\to\infty}\left(\frac{5}{3}\right)^{n}\sum_{j\sim_{n}i}g(u^{n}(v_{j})-u^{n}(v_{i}))\frac{(u^{n}(v_{j})-u^{n}(v_{i}))^{2}}{2}=0.$
We conclude that $\liminf{\mathcal{J}}_{n}(u^{n})\geq\mathcal{E}(u)$. This verifies the lower bound in the definition of $\Gamma$-convergence (cf. C1. Definition 3.1).
To verify the upper bound, we need to show that for an arbitrary $u\in X$ there is a recovery sequence $u^{n}\to u$ such that $\lim_{n\to\infty}{\mathcal{J}}_{n}(u^{n})={\mathcal{E}}(u^{\ast})$. To this end, let $u^{n}=u^{\ast}$ be a constant sequence. Arguing as above, we show that
$\lim_{n}{\mathcal{J}}_{n}(u)=\mathcal{E}(u).$ (3.7)
This completes the proof of $\Gamma$-convergence of ${\mathcal{J}}_{n}$.
Finally, the fact that convergent sequences of minimizers converge to a minimizer of the $\Gamma$-limit is a standard result (see, e.g., [2]). Below we include a short proof of this statement for completeness.
First, by $\Gamma$-convergence of $({\mathcal{J}}_{n})$, for $u\in X$ there is a recovery sequence $u^{n}\to u$ such that $\limsup_{n}{\mathcal{J}}_{n}(u^{n})\leq{\mathcal{E}}(u)$. Thus,
$\limsup_{n}\min_{w\in X}{\mathcal{J}}_{n}(w)\leq\limsup_{n}{\mathcal{J}}_{n}(u^{n})\leq{\mathcal{E}}(u).$
Since $u\in X$ is arbitrary,
$\limsup_{n}\min_{w\in X}{\mathcal{J}}_{n}(w)\leq\inf_{u\in X}\mathcal{E}(u).$ (3.8)
Suppose now that $u^{n}$ is a sequence of minimizers of ${\mathcal{J}}_{n}$ with $u^{n}\to u$. Using $\Gamma$-convergence of $({\mathcal{J}}_{n})$ again, we have
$\inf_{X}\mathcal{E}\leq\mathcal{E}(u)\leq\liminf_{n}{\mathcal{J}}_{n}(u^{n})=\liminf_{n}\min_{X}{\mathcal{J}}_{n}\leq\limsup_{n}\min_{X}{\mathcal{J}}_{n}\leq\inf_{X}{\mathcal{E}},$
where the final inequality follows from (3.8). Thus,
$\min_{X}{\mathcal{E}}={\mathcal{E}}(u)=\lim_{n}{\mathcal{J}}_{n}(u^{n}).$
The key implication of the $\Gamma$-convergence of $(\mathcal{E}_{n})$ and $(\mathcal{J}_{n})$ to $\mathcal{E}$ is that any limit point of a sequence of minimizers of $(\mathcal{E}_{n})$ or $(\mathcal{J}_{n})$ is a minimizer of $\mathcal{E}$, i.e., a harmonic function on the SG. For $(\mathcal{E}_{n})$, one can easily find the minimizers by either solving a system of linear equations or by harmonic extension. It is straightforward to show that the sequence of these minimizers converges to a harmonic function on SG. The situation for the Kuramoto energy functionals $(\mathcal{J}_{n})$ is more subtle: the minimization does not reduce to a linear system, and the harmonic extension method is not applicable. Nevertheless, using general properties of $\Gamma$-convergence, below we show the existence of a sequence of minimizers of $(\mathcal{J}_{n})$ that converges to a harmonic function on SG.
Theorem 3.4.
Let $u^{*}$ be the harmonic function on the SG subject to the following boundary condition $u^{*}|_{V_{0}}=\phi\in L(V_{0},\mathbb{R})$. Then there exists a sequence of stable equilibria of the KM on $\Gamma_{n}$, $u^{n}$, with $u^{n}|_{V^{0}}=\phi$ so that $\max_{x\in V_{n}}\|u^{n}(x)-u^{*}(x)\|\to 0$ as $n\to\infty$.
Proof.
Let $X$ be the space of Hölder continuous functions on $K$ with exponent $\beta=\frac{\log(5/3)}{2\log 2}$ equipped with the norm
$\|v\|_{\beta}\doteq\sup_{x\in K}|v(x)|+\sup_{x\neq y}\frac{|v(y)-v(x)|}{|x-y|^{\beta}}.$ (3.9)
Let $\varepsilon>0$ be arbitrary but fixed. Below, we prove that for all sufficiently large $n\in\mathbb{N}$, ${\mathcal{J}}_{n}$ has a local minimum in
$B_{\varepsilon}(u^{*})=\{f\in X\colon f|_{V_{0}}=\phi,\;\|f-u^{*}\|_{\beta}<\varepsilon\}.$
A continuous function on $K$, whose restriction to each $n$-cell of $K$ is a harmonic function is called a harmonic spline (cf. [18]). Consider the space of harmonic splines on $K$,
$X^{n}:=\{f\in C(K,\mathbb{R}):f\circ F_{w}\mbox{ is harmonic for all }|w|=n\},$ (3.10)
where $|w|$ stands for the length of the symbolic string $w=(w_{1},w_{2},\dots,w_{n})\in S^{n}$.
Then $X^{n}$ is a finite-dimensional subspace of $X$ and $X_{n}\cap\overline{B_{\varepsilon}(u^{*})}$ is a compact subset of $X$, as a closed and bounded subset of a finite-dimensional subspace. For every $n\in\mathbb{N}$, $J_{n}$ is a continuous functional on $X$. Thus, it achieves its minimum on $X_{n}\cap\overline{B_{\varepsilon}(u^{*})}$. Denote this minimum by $u^{n}$. In case of multiple minima, whenever possible we choose the one lying in $B_{\epsilon}(u^{*})$.
It remains to show that $u^{n}\in B_{\epsilon}(u^{*})$ for all but possibly finite number of $n\in\mathbb{N}$. That is, we show that for all $n$ sufficiently large there is a local minimizer, $u^{n}$ of ${\mathcal{J}}_{n}$ satisfying $\|u^{n}-u^{*}\|_{C(K)}<\varepsilon$, which will complete the proof.
Suppose on the contrary that there are infinitely many $n$ such that $\|u^{n}-u^{*}\|_{\beta}=\varepsilon$. Thus, there is an infinite subsequence $u^{n_{k}}\in C(K,\mathbb{R})$ such that $\|u^{n_{k}}-u^{*}\|_{\beta}=\varepsilon$ and ${\mathcal{J}}_{n}(u^{n_{k}})<{\mathcal{J}}_{n_{k}}(u^{*})$. The sequence $\{u^{n_{k}}\}$ is bounded in $X$ and using the compactness of the embedding $X$ in $C(K,\mathbb{R})$, we can extract a subsequence $u^{n_{j}}$ converging to $\bar{u}$ in the norm of $C(K,\mathbb{R})$.
By construction, the harmonic $u^{*}\in C(K,\mathbb{R})$ satisfies the following inequalities ${\mathcal{E}}_{n}(u^{*})\leq{\mathcal{E}}_{n}(\bar{u})\;\forall n\in\mathbb{N}.$ Therefore, ${\mathcal{E}}(u^{*})\leq{\mathcal{E}}_{n}(\bar{u})$ (cf. (2.8)). On the other hand, by Theorem 3.3,
$\mathcal{E}(\bar{u})\leq\liminf{\mathcal{J}}_{n}(u^{n_{j}})\leq\liminf{\mathcal{J}}_{n}(u^{*})=\mathcal{E}(u^{*}),$
where the first inequality follows from (3.2) and the second from (3.3). We conclude that $\mathcal{E}(\bar{u})=\mathcal{E}(u^{*})$ and $\bar{u}=u^{*}$. This contradicts the assumption that $\|u^{n}-u^{*}\|_{\beta}=\epsilon$. ∎
Remark 3.5.
Both Theorem 3.4 and Theorem 1.5 claim the existence of families of equilibria in the discrete KM models converging to harmonic functions on $K$. However, the settings of the two theorems differ. In the former case, we consider real-valued solutions subject to Dirichlet boundary conditions, whereas in the latter, we deal with $\mathbb{T}$-valued solutions for the model with free boundary conditions. Nonetheless, the proof of Theorem 3.4 provides a roadmap that will be useful in addressing the more challenging setting of Theorem 1.5.
4 The KM on nearest-neighbor graphs
In the previous section, we proved $\Gamma$-convergence of the Kuramoto energy to the Dirichlet energy as $n\to\infty$. Until now, we have treated solutions of the KM as real-valued functions. It is time to recognize the importance of viewing them as maps from their respective domains to the circle. To extend $\Gamma$-convergence techniques to $\mathbb{T}$-valued maps, we incorporate several geometric constructions developed in [14]. To make the transition to the $\mathbb{T}$-valued setting more intuitive, we first analyze the KM on the unit circle. The fact that a circle, an interval, and a cube are self-similar sets was used effectively in [18] to elucidate analysis on fractals. Similarly, we begin the analysis of the KM on SG by first considering it on nearest-neighbor graphs. Besides its pedagogical value, the KM on nearest-neighbor graphs is of independent interest, as it features prominently in many applications (see, e.g., [22, 1, 16]).
In this section, consider the KM on the nearest-neighbor graph
$\dot{u}^{n}(t,v_{i})=2^{n}\sum_{j\sim_{n}i}\sin\left(2\pi(u^{n}(t,v_{j})-u^{n}(t,v_{i}))\right),\quad i\in\left\{0,1,\dots,2^{n}-1\right\},$ (4.1)
where
$j\sim_{n}i\Leftrightarrow j=i\pm 1\mod 2^{n}$ (4.2)
Since we are interested in the continuum limit of (4.1) as $n\to\infty$, it is convenient to interpret $u^{n}(t,v_{i})$ as the values of $u^{n}(t,x)$ at $x=i2^{-n}$ where $i$ runs from 0 to $2^{n}-1$. Specifically, we represent the spatial domain of $u^{n}(t,x)$ as a discretization of the unit interval with the two end points identified $I=[0,1]\setminus\{0\simeq 1\}$ by the set of points
$V_{n}=\{0,2^{-n},2\cdot 2^{-n},\dots,(2^{n}-1)2^{-n}\}=:\{v_{0},v_{1},\dots\}.$ (4.3)
Accordingly, we view (4.1) as a dynamical system on the nearest neighbor graph $\Gamma^{n}$ with the vertex set $V_{n}$ (4.3) and the adjacency relation (4.2).
We will argue that the natural candidate for the continuum limit of (4.1) is the heat equation
$\partial_{t}u=\Delta u,$ (4.4)
on $(0,1)$ with periodic boundary condition $u(0)=u(1)$  [^2] In [8], the heat equation was identified as a continuum limit of the KM on geometric random graphs.. Here, we view $u(t,\cdot)$ as a map from $\mathbb{T}$ to $\mathbb{T}$. Thus, the steady states of (4.4) are harmonic maps from $\mathbb{T}$ to $\mathbb{T}$:
$\Delta u=0,\quad u(0)=u(1)=0.$ (4.5)
Here, we set $u(0)=0$ to eliminate translation invariance.
If $u$ were real-valued then $u\equiv 0$ would be a unique solution of (4.5). In the class of $\mathbb{T}$-valued functions, (4.5) has infinitely many solutions:
$u(x)=qx\mod 1,\quad q\in\mathbb{Z},$ (4.6)
i.e., one solution per homotopy class, determined by the degree $q$. Such solutions are called $q$-twisted states after [22].
Going back to the KM, it is easy to find discrete counterparts of (4.6):
$u(v_{i})=qi2^{-n}\mod 1,\quad i\in[2^{n}].$ (4.7)
In the case of the KM on the circle, (4.7) clearly converges to (4.6) as $n\to\infty$. However, for other self-similar sets, e.g., for SG, the relation between the solutions of the KM and those of its continuum limit is not as transparent. In preparation for the analysis of the KM on SG and other fractals, we extend below the $\Gamma$-convergence techniques from the previous section to cover $\mathbb{T}$-valued solutions.
Remark 4.1.
Below, we focus on nearest-neighbor ring graphs to make the connection with the KM on fractals, which is our ultimate goal. However, the analysis in the remainder of this section can be naturally extended to cover $k$-nearest-neighbor coupling. If $k\gtrsim O(1)$ the results can be further extended to random networks (e.g., small-world graphs) using approximation results for dynamical models on random networks [13]. In [8, 6], twisted states on geometric random graphs are analyzed using a different, albeit related, approach.
4.1 Harmonic maps on $\mathbb{T}$
Before we begin, we recall a few basic facts about continuous maps from $\mathbb{T}$ to $\mathbb{T}$ (see also Section 1). For a given $f\in C(\mathbb{T},\mathbb{T})$, there is a unique $\hat{f}\in C(\mathbb{R},\mathbb{R})$ such that
$\pi\circ\hat{f}(x)=f\circ\pi(x)\quad\forall x\in\mathbb{R},$
where $\pi(x)\doteq x\mod 1$. $\hat{f}$ is called the lift of $f$.
The integer $\hat{f}(1)-\hat{f}(0)$ is called the degree of $f$. The degree counts the number of rotations around the circle made by the trajectory $(f(t),\;0\leq t\leq 1)$. By the Hopf degree theorem, the degree determines the homotopy class of continuous maps from $\mathbb{T}$ to $\mathbb{T}$.
We now explain our strategy for extending the $\Gamma$-convergence techniques for the model at hand.
I.
Let $q\in\mathbb{Z}$ be fixed. For a homotopy class defined by $q$, we construct the covering of the circle. Define
$\mathbb{T}_{\times}$ $\doteq\mathbb{T}\times\mathbb{Z},$
$\mathbb{T}^{k}$ $\doteq\mathbb{T}\times\{k\}\subset\mathbb{T}_{\times},\;k\in\mathbb{Z}\qquad\mbox{(sheets)}$
Cut $\mathbb{T}^{k}$ at $0\times\{k\}$ producing two copies $0^{k}$ and $1^{k}$. Identify $1^{k}\simeq 0^{k+q}$. Then $\tilde{\mathbb{T}}\doteq\mathbb{T}_{\times}/\simeq$ is the covering space, and $\mathbb{T}^{0}$ is called the fundamental domain, see Figure 5. Discretize $\mathbb{T}^{0}$ by $V_{n}^{0}=\{0,2^{-n},\dots,(2^{n}-1)2^{-n},1\}$, where we omit superscripts on vertex points for notational ease.
Figure 5: Construction of the covering space $\tilde{\mathbb{T}}$ for $q=1$. The fundamental domain $\mathbb{T}^{0}$ is highlighted in red.
II.
On the fundamental domain, reformulate the boundary value problem for the discrete Laplacian for the lift of $u$:
$\Delta_{n}\hat{u}^{n}=0,\quad\hat{u}^{n}(0)=0,\;\hat{u}^{n}(1)=q.$ (4.8)
In one-dimension, this problem can be solved immediately:
$\hat{u}^{n}(v_{i})=qi2^{-n},\quad i\in\{0,1,\dots,2^{n}\}.$ (4.9)
III.
We have computed the value of solutions on a dense subset of the fundamental domain $V^{\ast}=\bigcup_{k\in\mathbb{Z}}V^{0}_{n}$. Noting that the discrete twisted states are (4.9) are Lipschitz continuous on $V^{\ast}$, we extend by them by continuity to the rest of the fundamental domain, to obtain the harmonic function on the fundamental domain, denoted $\hat{u}^{\ast}.$ Extend these solutions to other sheets of the covering space:
${\bf\hat{u}^{\ast}}(x,k)=\hat{u}^{\ast}(x)+k,\quad k\in\mathbb{Z}.$
Then $^{\bf\hat{u}^{\ast}}$ is a harmonic function on the covering space.
IV.
Finally, project the range of the harmonic function $^{\bf\hat{u}^{\ast}}$ back to $\mathbb{T}$-values:
$u^{*}:={\bf\hat{u}^{\ast}}|_{\mathbb{T}^{0}}\mod 1.$
Then $u^{*}$ is the desired harmonic map on $\mathbb{T}$.
4.2 $\Gamma$-convergence on the covering space
In Section 3, we demonstrated convergence of the energy functional in the KM to the Dirichlet energy, from which we derived convergence of stable equilibria of the KM to harmonic functions. To derive these statements to the KM on the nearest-neighbor graph, we first need to translate the analysis to the real-valued setting. To this end, we use the procedure described in the previous subsection. Specifically, we first lift the $\mathbb{T}$-valued solutions on $\mathbb{T}$ and their discretizations to $\mathbb{R}$-valued functions on the covering spaces. Then we replicate the arguments from Section 3 to obtain desired convergence of the energy functionals and the equilibria on the fundamental domain, $\mathbb{T}^{0}$.
Discrete harmonic functions $\hat{u}^{n}$ minimize
${\mathcal{E}}_{n}({u}):=2^{n}\sum_{(i,j)\in E_{n}}\frac{({u}(v_{j})-{u}(v_{i}))^{2}}{2},\quad{u}=({u}(v_{0}),{u}(v_{1}),\dots,{u}(v_{2^{n}})).$
over $X_{q}^{n}=\{f\in\mathbb{R}^{2^{n}+1}:\;f(v_{0})=0,\;f(v_{2^{n}})=q\}$, the discrete approximation of the fundamental domain $\mathbb{T}^{0}$ with fixed jump condition.
Let $X_{q}$ be the space
$X_{q}=\{f\in C([0,1],\mathbb{R}):f(0)=0,\;f(1)=q\}.$
As in (2.8), the limit
${\mathcal{E}}(f)=\lim_{n\to\infty}{\mathcal{E}}_{n}(f)$ (4.10)
is well-defined. The domain of $\mathcal{E}$}, $\operatorname{dom}({\mathcal{E}})=H^{1}([0,1])$ [18].
To apply the techniques from the proof of Theorem 3.3, we need only establish Hölder continuity of functions in the domain of $\mathcal{E}$}, cf. Lemma 2.2. By Morrey’s inequality [4], we have
$\|{u}\|_{1/2}\leq C\|{u}\|_{H^{1}([0,1])},$
where $\|\cdot\|_{1/2}$ is the $1/2$-Hölder norm, and $C$ is independent of $u$. Thus,
$\frac{|{u}(x)-{u}(y)|}{|x-y|^{1/2}}\leq\|{u}\|_{1/2}\leq C_{1}\|{u}\|_{H^{1}([0,1])}\leq C_{2}\|{u}^{\prime}\|_{L^{2}([0,1])}=C_{2}\sqrt{\mathcal{E}({u})}.$
Here, we also used the Poincare inequality to bound $\|{u}\|_{H^{1}([0,1])}$.
Now we are in a position to use the arguments from Section 3. Fix the space $X_{q}$ with the supremum metric. As before, $\Gamma-\lim{\mathcal{E}}_{n}={\mathcal{E}}$ due to the monotone pointwise convergence ${\mathcal{E}}_{n}\nearrow{\mathcal{E}}$ and continuity of the ${\mathcal{E}}_{n}$.
Next, we define the Kuramoto energy on on the fundamental domain $\mathbb{T}^{0}$
${\mathcal{J}}_{n}({u}):=2^{n}\cdot\frac{1}{4\pi^{2}}\sum_{(i,j)\in E_{n}}(1-\cos(2\pi({u}(v_{j})-{u}(v_{i})))).$ (4.11)
Proceeding as in Theorem 3.3, we obtain $\Gamma$-convergence for (real-valued) energies on the fundamental domain.
Theorem 4.2.
Let $({\mathcal{J}}_{n})$ be the Kuramoto energy functionals (4.11) and let $\mathcal{E}$} be given by (4.10). Then
${\mathcal{E}}=\Gamma\dash\lim_{n\to\infty}{\mathcal{J}}_{n}.$ (4.12)
Moreover, suppose $({u}^{n})$ is a sequence of minimizers of $({\mathcal{J}}_{n})$ converging to ${u}^{\ast}\in X_{q}$. Then $u^{\ast}$ is a minimizer of $\mathcal{E}$ and
${\mathcal{E}}({u}^{\ast})=\lim_{n}{\mathcal{J}}_{n}({u}^{n}).$ (4.13)
Remark 4.3.
Theorem 4.2 establishes convergence of (real-valued) minimizers on the covering space $\mathbb{T}^{0}$. However, by periodicity of $\sin$ and $\cos$, it is clear that the minimizers ${u}^{n}$ remain minimizers of the associated $\mathbb{T}$-valued energy when projected to harmonic maps ${u}^{n}\mod 1$.
The analogue of Theorem 3.4 is trivial in this setting. Indeed, it is known that for all $n$ sufficiently large (depending on $q$), the discrete twisted states (4.7) are stable equilibria of the KM. Thus, the minimizers of ${\mathcal{E}}_{n}$ and ${\mathcal{J}}_{n}$ coincide for sufficiently large $n$.
Remark 4.4.
It is interesting to relate Theorem 4.2 to the stability analysis of equilibria in the KM presented in [1]. In addition to the twisted states, whose stability can be established using the discrete Fourier transform (cf. [15, Theorem 3.4]), there exists another class of equilibria – half-twisted states:
${u}^{(r,n)}(v_{i})=\frac{r}{2^{n}-2}i,\qquad i\in\{1,\dots,2^{n}\},$
where $r\in\mathbb{Z}+1/2$ is a half-integer satisfying $\frac{2^{n}}{4}+\frac{1}{2}<r<\frac{2^{n}}{4}-\frac{1}{2}$ (see Fig. 6).
Denote the limit of ${u}^{(r,n)}$ by ${u}^{r}$. If the sequence $({u}^{(r,n)})$ were composed of minimizers of the Kuramoto energy ${\mathcal{J}}_{n}$, then by Theorem 4.2, the limit $u^{r}$ would be a minimizer of $\mathcal{E}$}. However, since $u^{r}$ is discontinuous, it lies outside $\operatorname{dom}({\mathcal{E}})$ . We therefore conclude that ${u}^{(r,n)}$ cannot be minimizers of ${\mathcal{J}}_{n}$ for sufficiently large $n$. This is consistent with the results of [1], which show that ${u}^{(r,n)}$ are saddles.
Figure 6: A half-twisted state $u^{(r,n)}$ as described in Remark 4.4. Here $r=1/2$ and $n=3$.
5 Harmonic maps from SG to the circle
Extending the ideas of the previous section, we aim to show that the stable equilibria of the KM on graphs approximating SG converge to minimizers of the Dirichlet energy for $\mathbb{T}$-valued functions on SG, i.e., harmonic maps. As before, our approach relies on $\Gamma$-convergence techniques, now adapted to the setting of $\mathbb{T}$-valued functions on SG. To this end, we construct an appropriate covering space for SG and lift $\mathbb{T}$-valued functions on SG to real-valued functions on this covering space. We then apply $\Gamma$-convergence to show that the stable equilibria in the KM are close to harmonic maps, which minimize the Dirichlet energy on the covering space – more precisely, on a fundamental domain. Finally, we project back to SG to establish convergence of the equilibria of the KM to harmonic maps.
In Sections 5.1 and 5.2, we explain the construction of the covering space and the definition of Dirichlet energy for the problem at hand. This material, adapted from [14], is included for completeness. As in the nearest-neighbor graph setting, the covering space depends on the degree of the harmonic map and is constructed separately for each homotopy class. For simplicity, Sections 5.1 and 5.2 focus on simple harmonic maps, whose degrees are given by a single integer. In Section 5.3, we extend the discussion to higher-order maps, where degrees are represented by vectors.
After that we prove the convergence of stable equilibria in the KM to a unique harmonic map within the corresponding homotopy class. This is the goal of Section 6, where we follow the approach of Section 3 to prove Theorem 1.5, the main result of this paper.
5.1 The covering space
As in Section 4, the key step in the construction of harmonic maps is setting up the covering space for $K$. This space is constructed separately for each degree vector $\bar{\omega}(u)$. To avoid obscuring the main ideas behind the construction with technical details required for the general case, we focus in this subsection on the special case where $\bar{\omega}(u)=(\rho_{0}),\;\rho_{0}\neq 0$, i.e., when the winding number along the boundary of the base triangle $T$ is nontrivial and equal to $\rho_{0}$, while the winding numbers along all other loops $\partial T_{w},|w|>0$ are trivial. In the following subsection, we explain the additional details needed to handle the general case.
As in Section 4.1, the construction of the covering space for $K$ consists of three steps: stack, cut, and identify (see Figure 7):
  1. 1.
    First, let
    $K_{\times}$ $\doteq K\times\mathbb{Z},$ (5.1)
    $K^{s}$ $\doteq K\times\{s\}\subset K_{\times},\quad s\in\mathbb{Z},$ (5.2)
    $K^{s}_{i}$ $\doteq F_{i}(K^{s}),\quad i\in S,\quad s\in\mathbb{Z},$ (5.3)
    $K^{s}_{1}\cap K^{s}_{2}=\{x^{s}\},$ $\quad K_{2}^{s}\cap K_{3}^{s}=\{y^{s}\},\quad K_{3}^{s}\cap K_{1}^{s}=\{z^{s}\},\quad s\in\mathbb{Z},.$ (5.4)
    Figure 7: The construction of the covering space of $K$ corresponding to the degree $\bar{\omega}(u)=(1)$.
  2. 2.
    Cut each $K^{s}$ at $z^{s}$, i.e., replace $z^{s}$ with two distinct copies
    $z^{s}_{-}=v^{s}_{1\bar{3}}\quad\mbox{and}\quad z^{s}_{+}=v^{s}_{3\bar{1}}.$ (5.5)
  3. 3.
    Identify
    $z^{s}_{+}=v_{3\bar{1}}^{s}\simeq v_{1\bar{3}}^{s+\rho_{0}}=z^{s+\rho_{0}}_{-},\qquad s\in\mathbb{Z}.$ (5.6)
    The resultant covering space is then $\tilde{K}\doteq K_{\times}/\simeq$. The copies of $K$, belonging to different levels $s\in\mathbb{Z}$, compose the sheets of $\tilde{K}$. We keep denoting them by $K^{s}$. Each sheet contains both copies of $z^{s}:$ $z_{-}^{s}$ and $z^{s}_{+}.$ $K^{0}$ is called the fundamental domain.
Finally, we introduce a family of graphs $\Gamma_{m}^{s}=(V_{m}^{s},E_{m}^{s})$ approximating $K^{s}$. The $\Gamma_{m}^{s}$ are constructed in the same way as the approximating graphs $\Gamma_{m}$ of SG (see Figure 2), with the only distinction that $z^{s}$ is replaced with the two copies $z^{s}_{-}$ and $z^{s}_{+}$ (see Figure 8). By identifying $V_{m}^{s}\ni z_{+}^{s}\simeq z_{-}^{s+\rho_{0}}\in V_{m}^{s+\rho_{0}}$, we obtain the discretization of the covering space $\tilde{K}$, denoted $\tilde{\Gamma}_{m},\;m\in\mathbb{N}$. The set of nodes of $\tilde{\Gamma}_{m}$ is $\tilde{V}_{m}$. Then $\tilde{V}_{\ast}=\bigcup_{m=0}^{\infty}\tilde{V}_{m}$, and $\tilde{V}_{\ast}^{s}=\bigcup_{m=1}^{\infty}V_{m}^{s}.$ As before, $\tilde{V}_{\ast}$ is dense in $\tilde{K}$.
Figure 8: Approximating graphs $\Gamma_{m}^{k}$ for each sheet $K^{s}$ of the covering space $\tilde{K}$. Since our construction results in two copies of the vertex $z$, the resultant graphs are distinct from Figure 2. Superscripts $s$ are suppressed for simplicity.
5.2 Harmonic structure on $\tilde{K}$
In this subsection, we construct a harmonic map from $K$ to $\mathbb{T}$ without imposing any boundary conditions. To this end, we first construct a real-valued harmonic function on the fundamental domain $K^{0}$, which will extend to the covering space $\tilde{K}$.
The energy form on $\Gamma^{0}_{m}$ is defined as follows
${\mathcal{E}}_{m}[\Gamma^{0}_{m}](f)=\left(\frac{5}{3}\right)^{m}\sum_{(x,y)\in E_{m}^{0}}\frac{\left(f(x)-f(y)\right)^{2}}{2},\qquad f\in L(V_{m}^{0},\mathbb{R}),m\in\mathbb{N}.$ (5.1)
We consider minimizing ${\mathcal{E}}_{m}[\Gamma^{0}_{m}](f)$ subject to the jump condition
$f(z^{0}_{+})=f(z^{0}_{-})+\rho_{0},$ (5.2)
As before, to remove translation invariance we fix $f(v^{0}_{1})=0$. Since we will work exclusively on the fundamental domain $K^{0}$, in the following we suppress some superscripts for simplicity.
The minimization problem can be reformulated as
${\mathcal{E}}_{m}[\Gamma^{0}_{m}](f)\longrightarrow\;\min_{f\in H^{0}_{m}},$ (5.3)
where
$H^{0}_{m}\doteq\{f\in L(V_{m}^{0},\mathbb{R}):\quad f(v_{1})=0,\;f(z_{+})=f(z_{-})+\rho_{0}\}.$ (5.4)
The energy form ${\mathcal{E}}_{m}[\Gamma^{0}_{m}]$ inherits the key properties of the Dirichlet form ${\mathcal{E}}_{m}$ (see (2.5), (2.6)):
  1. 1.
    A $\Gamma_{m}^{0}$-harmonic [^3] In the present context, a function $f$ is said to be $\Gamma_{m}^{0}$-harmonic if it satisfies the discrete Laplace equation: $\sum_{(x,y)\in E_{m}^{0}}(f(y)-f(x))=0,\qquad\forall x\in V_{m}^{0}\setminus\{v_{1},z_{+},z_{-}\}.$ (5.5) $f_{m}^{0}\in H_{m}^{0}$ minimizes ${\mathcal{E}}_{m}[\Gamma^{0}_{m}]$ over $H_{m}^{0}$:
    ${\mathcal{E}}_{m}[\Gamma^{0}_{m}](f^{0}_{m})=\min\left\{{\mathcal{E}}_{m}[\Gamma^{0}_{m}](f):\;f\in H_{m}^{0}\right\}.$ (5.6)
  2. 2.
    The minimum of the energy form ${\mathcal{E}}_{m}[\Gamma^{0}_{m}]$ over all extensions of $f\in H_{m-1}^{0}$ to $H_{m}^{0}$ is equal to ${\mathcal{E}}_{m}[\Gamma^{0}_{m-1}](f)$:
    $\min\left\{{\mathcal{E}}_{m}[\Gamma^{0}_{m}](\tilde{f}):\;\tilde{f}\in H_{m}^{0},\;\tilde{f}|_{V_{m-1}}=f\in H_{m-1}^{0}\right\}={\mathcal{E}}_{m}[\Gamma^{0}_{m-1}](f).$ (5.7)
    Moreover, this minimizer is obtained by taking the harmonic extension of $f$.
As in (2.8), since energies are monotone increasing, the limit
${\mathcal{E}}_{K^{0}}(f)=\lim_{m\to\infty}{\mathcal{E}}_{m}[\Gamma_{m}^{0}](f)$ (5.8)
is well defined for $f\in C(K^{0},\mathbb{R})$.
Applying the arguments in [14, Lemma 5.1], we also conclude that the variational problem (5.3) has a unique solution $f^{0}_{m}\in H^{0}_{m}$:
${\mathcal{E}}_{m}[\Gamma_{m}^{0}](f^{0}_{m})=\min_{f\in H_{m}^{0}}{\mathcal{E}}_{m}[\Gamma_{m}^{0}](f),$
and the minimizers $f^{0}_{m}$ satisfy the following properties:
  1. 1.
    $f_{m}^{0}$   is $\Gamma^{0}_{m}$-harmonic,
  2. 2.
    $f^{0}_{m+1}|_{V^{0}_{m}}=f_{m}^{0}$.
Remark 5.1.
[…]
Figure 9: The boundary conditions for the minimization problem (5.3): the values of the function is prescribed only at $v_{1}$ and the jump condition is imposed at $z_{-}$ and $z_{+}$.
Thus, minimizers of (5.3) are constructed by first minimizing ${\mathcal{E}}_{1}[\Gamma_{1}^{0}]$, followed by repeated application of the harmonic extension. Let $\hat{f}_{m}$ be the minimizer of $E[\Gamma_{m}^{0}](f)$. By (5.7), ${\mathcal{E}}_{m}[\Gamma_{m}^{0}](\hat{f}_{m})={\mathcal{E}}_{1}[\Gamma_{1}^{0}](\hat{f}_{1})$ and $\hat{f}_{m}|_{V_{1}}=\hat{f}_{1}$.
So, we first find the values of $\hat{f}_{1}$ at points $x,y,z:=z_{+},v_{2},v_{3}$, by minimizing ${\mathcal{E}}_{1}[\Gamma_{1}^{0}]$ (see Fig. 9). Direct calculation yields the unique solution
$\begin{pmatrix}\hat{f}_{1}(x)\\ \hat{f}_{1}(v_{2})\\ \hat{f}_{1}(y)\\ \hat{f}_{1}(v_{3})\\ \hat{f}_{1}(z)\end{pmatrix}=\begin{pmatrix}1\\ 1\\ 1\\ 1\\ 1\end{pmatrix}\hat{f}_{1}(v_{1})+\begin{pmatrix}1/6\\ 2/6\\ 3/6\\ 4/6\\ 5/6\end{pmatrix}\rho_{0}.$
Using the harmonic extension (2.9) in the subdomains $v_{1}xz_{-}$, $xv_{2}y$, and $z_{+}yv_{3}$ inductively, we obtain a minimizer $\hat{f}^{*}$, on $V_{*}^{0}$.
We extend it by continuity to the fundamental domain $K^{0}$, and further to a uniformly continuous function on the covering space:
$\mathbf{\hat{f}}^{\ast}(x,s)=\hat{f}^{\ast}(x)+s,\qquad x\in K^{0},\;s\in\mathbb{Z}.$ (5.9)
Recall that $\tilde{\Gamma}_{m}$ and $\tilde{V}_{m}$ stand for the family of graphs approximating the covering space $\tilde{K}$ and their vertex sets respectively (see item $5$ in §5.1). By [14, Theorem 5.2], the restriction of $\mathbf{\hat{f}}^{\ast}$ on $\tilde{V}_{m}$, $\mathbf{\hat{f}}^{\ast}|_{\tilde{V}_{m}}$, is $\tilde{\Gamma}_{m}$-harmonic for every $m\in\mathbb{N}$, i.e., it satisfies the discrete Laplace equation
$\Delta_{m}\mathbf{\hat{f}}^{\ast}|_{\tilde{V}_{m}}(x)=0\quad\forall x\in\tilde{V}_{m}\setminus\{v^{s}_{1}\}.$ (5.10)
Finally, by restricting the domain of $\mathbf{\hat{f}}^{\ast}$ to the fundamental domain and by projecting the range of $\mathbf{\hat{f}}^{\ast}$ to $\mathbb{T}$, we obtain the harmonic map from the SG to the circle $f^{\ast}:\mskip 3.0muK\to\mathbb{T}$:
$f^{\ast}\doteq\mathbf{\hat{f}}^{\ast}|_{K^{0}}\mod 1.$ (5.11)
Hölder continuity of functions on $K^{0}$ also holds with minor adjustments. The fundamental domain $K^{0}$ is the union of the three subdomains $K^{0}_{i},\;i=1,2,3$ (see Fig. 8), each a copy of $K$ itself. Specifically, $K^{0}_{1},K^{0}_{2},$ and $K^{0}_{3}$ correspond to the intersection of $K^{0}$ with the closed triangles $v_{1}xz_{-},$ $xv_{2}y$, and $z_{+}yv_{3}$, respectively. Let $f\in\operatorname{dom}({\mathcal{E}}_{K^{0}})$. By Lemma 2.2, we immediately see that the restriction of $f$ on each of the subdomains $K^{0}_{i},i\in[3]$ is a Hölder continuous function. This yields the following lemma.
Lemma 5.2.
Let $f\in\operatorname{dom}({\mathcal{E}}_{K^{0}})$ then for every $K_{i}^{0},i\in[3]$ and $x,y\in K^{0}_{3}$,
$\frac{|f(x)-f(y)|}{|x-y|^{\beta}}\leq C\sqrt{{\mathcal{E}}_{K^{0}}(f)},\quad\beta=\frac{\log(5/3)}{2\log(2)},$
where $C>0$ is independent of $f$.
We conclude this subsection by showing that minimization of the Dirichlet energy on the fundamental domain without imposing any boundary conditions on $V_{0}$ results in a solution satisfying zero Neumann boundary conditions.
Lemma 5.3.
$f^{\ast}$ satisfies natural boundary conditions (1.14).
Proof.
We recall that for a harmonic function $\hat{f}^{*}\in C(K^{0},\mathbb{R})$, the normal derivative at a boundary vertex $v_{j}\in V_{0}$ is
$\partial_{\bf n}\hat{f}^{*}(v_{j})=\left(\frac{5}{3}\right)^{m}\sum_{y\sim_{m}j}\hat{f}^{*}(y)-\hat{f}^{*}(v_{j}),$
which is well-defined because the right hand side is constant in $m$ for harmonic functions [18]. In the following, we work always on the fundamental domain and suppress superscripts/subscripts for simplicity.
For every $m\in\mathbb{N}$, the Euler-Lagrange equations for (5.1) at the boundary points $v_{2}$ and $v_{3}$ is
$\sum_{j\sim_{m}i}(\hat{f}^{*}(v_{j})-\hat{f}^{*}(v_{i}))=0,\quad i=2,3,$ (5.12)
from which it follows $\partial_{\bf n}\hat{f}^{*}(v_{2})=\partial_{\bf n}\hat{f}^{*}(v_{3})=0$. To see that $\partial_{\bf n}\hat{f}^{*}(v_{1})=0$ as well, we use a discrete divergence theorem. For $x\in\Gamma^{0}_{n}\setminus V^{0}_{0}$, define
[…]
This way, $\Delta_{n}\hat{f}^{*}(x)=0$ for all $x\not\in V_{0}$ (cf. (5.10)). Then, summing over all junction points results in the discrete divergence theorem:
$\sum_{x\in V_{n}\setminus V_{0}}\Delta_{n}\hat{f}^{*}(x)=\sum_{v_{i}\in V_{0}}\sum_{y\sim_{n}v_{i}}\hat{f}^{*}(y)-\hat{f}^{*}(v_{i}).$
Together with (5.12) this shows $\partial_{\bf n}\hat{f}^{*}(v_{1})=0$. Projecting $f^{*}=\hat{f}^{*}|_{\mathbb{T}}$ preserves normal derivatives, so the result follows. ∎
5.3 Higher order harmonic maps
In the previous section, we explained how to construct a harmonic map when the winding vector has a single nontrivial entry, $\bar{\omega}(f)=(\rho_{0})$. The procedure for a general winding vector, $\bar{\omega}(f)=(\rho_{0},\rho_{1},\dots,\rho_{k})$, follows the same general approach, but it requires additional cuts to construct the covering space. The idea is to introduce one cut point for each nonzero entry in the winding vector. Care must be taken in choosing these points, as the loops used for cutting must be linearly independent. The choices of cut points when $\bar{\omega}(f)=(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ and $\bar{\omega}(f)=(\rho_{0},\rho_{1},\rho_{2},\dots,\rho_{12})$ are shown in Figure 10. The general procedure of selecting cut points for a p.c.f. fractal is explained in detail in Sections 7 and 8 of [14]. For completeness, we discuss the construction of the covering space for SG for an arbitrary winding vector below and refer the reader to [14] for further details.
Let $f\in C(K,\mathbb{T})$ be a harmonic map of order $N\in\mathbb{N}$, i.e., $\bar{\omega}(f)=(\rho_{0},\rho_{1},\dots,\rho_{k})$ with
$\frac{3^{N}-1}{2}\leq k\leq\frac{3^{N+1}-3}{2}.$
Note that to each nonzero entry of $\bar{\omega}(f)$, $\rho_{i},i\leq\frac{3^{N+1}-3}{2}$, corresponds a unique loop $\partial T_{w},w\in S^{N}$. For the remainder of this subsection, let us denote $\rho_{i}$ by $\rho_{w}$, i.e., by the itinerary of the cell, whose boundary has winding number $\rho_{i}$.
For each nonzero $\rho_{w}$ with $n:=|w|\leq N$, we cut $K^{s}$ at a vertex $z_{w}$ satisfying $z_{w}\in\partial T_{w}\setminus(\bigcup_{|v|=n-1}\partial T_{v}),$ see Fig. 10. Denote the two copies of $z_{w}$ by $(z_{w})_{-}$ and $(z_{w})_{+}$. Finally, identify vertices in the sheets $K^{s}$ by
$(z_{w}^{k})_{+}\simeq(z_{w}^{k+\rho_{w}})_{-}.$
The covering space is then $\tilde{K}\doteq K_{\times}/\simeq$, and the fundamental domain $K^{0}$. Construction of the approximating graphs of the covering space, $\Gamma^{s}_{m}=(V^{s}_{m},E^{s}_{n})$, follows analogously.
Figure 10: For arbitrary degree $\bar{\omega}(f)$, each non-zero entry requires an associated cut vertex. As before, for $\partial T$ we cut at the single orange vertex. For loops $\partial T_{1},\partial T_{2},\partial T_{3}$ we cut the three blue vertices. Finally, for loops $\partial T_{ij}$ there are 9 corresponding green vertices.
Constructing the harmonic structure on $\tilde{K}$ also follows along the same lines as before. We ntroduce the corresponding spaces on the fundamental domain $K^{0}$:
$H^{0}_{m}=\{f\in L(V^{0}_{m},\mathbb{R})\colon$ $f(v^{0}_{1})=0,$
$\;f((z^{0}_{w})_{+})=f((z^{0}_{w})_{-})+\rho_{w},\forall w\in S^{n},|n|\leq m\}$
and energies ${\mathcal{E}}_{m}[\Gamma_{m}^{0}]$ as in (5.1). The resultant harmonic structure satisfies the same results as in Section 5.2, so we construct a harmonic function $^{\bf\hat{f}}^{*}$ on the covering space by minimizing the ${\mathcal{E}}_{m}[\Gamma_{m}^{0}]$ on the fundamental domain within the class $H_{m}^{0}$, followed by inductive application of the harmonic extension algorithm. The projection of the range of $^{\bf\hat{f}}^{*}$ to $\mathbb{T}$ yields the desired HM, $f^{*}$, with degree $\bar{\omega}$.
Arguing as in § 5.2, we represent the fundamental domain as a union of closed $m$-cells $G^{0}=\bigcup_{|w|=m}K^{0}_{w}$. Recall $K^{0}_{w}=F_{w}(K^{0})$. As before, the restriction of $f\in\operatorname{dom}({\mathcal{E}}_{K^{0}})$ on each of these subdomains is a Hölder continuous function.
Lemma 5.4.
Let $f\in\operatorname{dom}({\mathcal{E}}_{K^{0}})$ and $K^{0}_{w},$ $|w|=m$. For any $x,y\in K^{0}_{w}$ we have
$\frac{|f(x)-f(y)|}{|x-y|^{\beta}}\leq C\sqrt{{\mathcal{E}}_{K^{0}}(f)},\quad\beta=\frac{\log(5/3)}{2\log(2)},$
where $C$ is independent of $f$.
6 The proof of Theorem 1.5
We are now prepared to prove Theorem 1.5. We work with the lifts of solutions of the KM restricted to the fundamental domain. The proof for the most part coincides with the proofs of Theorems 3.3 and 3.4. Only minor adjustments are required to deal with the fact that the covering space is a union of finitely many copies of $K$. Below, we comment on the necessary adjustments to the arguments in Section 3.
Let
$\omega^{\ast}=\left(\rho_{0},\rho_{1},\dots,\rho_{M(m)}\right)\in\mathbb{Z}^{M(m)+1},M(m)\doteq\frac{3(3^{m}-1)}{2},$
be arbitrary but fixed. We assume that
$\sum_{i=M(m-1)+1}^{M(m)}|\rho_{i}|>0.$
The last condition means that at least some cells of order $m$ are used in the construction of the covering space corresponding to $\omega^{\ast}$.
Next, construct the covering space using the algorithm explained in §§5.2, 5.3. Decompose the fundamental domain $K^{0}$ into the union of $m$-cells, $K^{0}=\bigcup_{|w|=m}K^{0}_{w}$. Let $\hat{u}_{m}$ stand for the unique minimizer of the Dirichlet energy ${\mathcal{E}}_{m}[\Gamma_{m}^{0}]$ over $H^{0}_{m}$ and denote by $\hat{u}^{*}$ the harmonic extension of $\hat{u}_{m}$ to $K^{0}$. Finally, $u^{\ast}$ is obtained by projecting the co-domain of $\hat{u}^{*}$ to $\mathbb{T}$. By construction (and using Lemma 5.3), $u^{\ast}$ satisfies (1.12) - (1.15).
$\Gamma$-convergence. We first establish $\Gamma$-convergence of the Kuramoto energies on $K^{0}$. Take as $X$ the space of continuous functions $C(K^{0},\mathbb{R})$ with the supremum metric. As before, $\Gamma\dash\lim_{n}{\mathcal{E}}_{n}={\mathcal{E}}_{K^{0}}$ by virtue of continuity of the ${\mathcal{E}}_{n}$ and monotone pointwise convergence of the energies. Next define the Kuramoto energy on $\Gamma_{n}^{0}$:
${\mathcal{J}}_{n}(u)=\left(\frac{5}{3}\right)^{n}\frac{1}{4\pi^{2}}\sum_{(i,j)\in E_{n}^{0}}(1-\cos(2\pi(u(v_{j})-u(v_{i})))).$ (6.1)
Repeating the proof of Theorem 3.3 establishes $\Gamma\dash\lim_{n}{\mathcal{J}}_{n}={\mathcal{E}}_{K^{0}}$. We remark only that when establishing estimates (3.6), each pair of adjacent vertices live in a common $m$-cell so that Hölder estimates of Lemma 5.4 apply.
Existence of sequence of stable equilibria. It remains to show that there is a sequence of stable steady states of the KM, $u^{n}\in L(V_{n}^{0},\mathbb{T}),$ converging to $u^{\ast}$.
We proceed as in Theorem 3.4 to construct minimizers of ${\mathcal{J}}_{n}$ converging to $u^{*}$. Importantly, since we now seek minimizers in the space $H_{m}^{0}$ (which does not assume Dirichlet boundary conditions), then minimizers of ${\mathcal{J}}_{n}$ correspond to stable equilibria of (1.3) (cf. (1.6)). Since the proofs are essentially the same, we only outline the main steps.
Recall that $K^{0}=\bigcup_{|w|=m}K^{0}_{w}$. Introduce the space of Hölder continuous functions on each $K^{0}_{w}$ with norm
$\|f\|_{C^{\beta}(K^{0}_{w})}=\sup_{x\in K^{0}_{w}}|f(x)|+\sup_{x\neq y}\frac{|f(y)-f(x)|}{|x-y|^{\beta}}.$
Let $X$ be the space of functions on $K^{0}$ that are Hölder continuous on each $K^{0}_{w}$ with exponent $\beta=\frac{\log(5/3)}{2\log 2}$. Fix $\varepsilon>0$ and consider the minimization of ${\mathcal{J}}_{n}$ on the closed set
$\overline{B_{\varepsilon}(u^{*})}:=\{f\in X\colon\max_{w}\|f-u^{*}\|_{C^{\beta}(K^{0}_{w})}\leq\varepsilon\}.$
Restricting the minimization problem to the space of harmonic splines, $X^{n}$, (cf. (3.10)) is a finite dimensional problem, and so repeating the compactness argument of ${\mathcal{J}}_{n}$ guarantees a minimum on $X^{n}\cap\overline{B_{\varepsilon}(u^{*})}$.
We need only show that minimizers are attained on the interior of the ball $B_{\varepsilon}(u^{*})$ for all $n$ sufficiently large (so as to be true local minima). Suppose not. Then, there is a sequence $f^{n}\in X$ with $\max_{w}\|f^{n}-u^{*}\|_{C^{\beta}(K^{0}_{w})}=\varepsilon$. Restricting to $K^{0}_{1}$, Lemma 5.4 implies that there is a subsequence $f^{n_{k}}\to\bar{f}\in C(K^{0}_{1})$. If necessary, pass to a further subsequence so that $f^{n_{k_{\ell}}}\to\bar{f}\in C(K^{0}_{1}\cup K^{0}_{2})$. As there are finitely many subdomains $K^{0}_{w}$, repeating this results in a subsequence (which we denote $f^{n_{j}}$ for simplicity) that converges on the entire domain: $f^{n_{j}}\to\bar{f}\in C(K^{0})$.
As before, by the $\Gamma$-convergence of ${\mathcal{J}}_{n}$ to ${\mathcal{E}}_{K^{0}}$, we obtain the contradiction $\bar{f}=u^{*}$. Thus, for all $n$ sufficiently large, there is a (local) minimizer, $u^{n}$, of ${\mathcal{J}}_{n}$ in the interior of $B_{\varepsilon}(u^{*})$. By construction $u^{n}$ has degree $\omega^{\ast}$. Thus both (1.16) and (1.17) follow.
7 The KM on post-critically finite fractals
The analysis of functions on SG naturally extends to a broader class of self-similar domains known as p.c.f. fractals. Examples include higher-order Sierpinski Gaskets, the pentagasket, and the hexagasket (see Fig. 11), to name a few. In this final section, we describe how the preceding results can be extended to the KM on p.c.f. fractals. While there are some technical differences in the construction of covering spaces (highlighted below), the techniques developed above remain applicable.
Figure 11: Representative examples of p.c.f. fractals: the three-level SG, $SG_{3}$, the pentagasket, and the hexagasket. For details on their construction see [18, 14].
7.1 Background on p.c.f. fractals
We first review some relevant background of p.c.f. fractals and refer the reader to [10] for further details. Let $K\subset\mathbb{R}^{d}$ be a non-empty, connected, and compact set satisfying
$K=\bigcup_{i=1}^{N}F_{i}(K),$ (7.1)
where the $F_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $1\leq i\leq N$, are contraction maps with corresponding contraction ratios $0<\lambda_{i}<1$. For simplicity, assume the $F_{i}$ are homotheties so that $|F_{i}(x)-F_{i}(y)|=\lambda_{i}|x-y|$. Denoting the $n$-cells
$K_{w}=F_{w}(K),\quad w=(w_{1},w_{2},\dots,w_{n})\in[N]^{n},$
the critical set is defined by
$\mathcal{C}=\bigcup_{i\neq j}(K_{i}\cap K_{j}).$
If $\mathcal{C}\neq\emptyset$, then the post-critical set is
$V_{0}=\bigcup_{m\geq 1}\bigcup_{|w|=m}F_{w}^{-1}(\mathcal{C}).$
If $V_{0}$ is finite, then $K$ is called a p.c.f. fractal. Moreover, the post-critical set $V_{0}$ is taken to be the boundary of $K$ (e.g., for the associated Dirichlet problem).
We assume there is a harmonic structure defined on $K$, which we describe below. Define the vertex sets
$V_{n}=\bigcup_{i=1}^{N}F_{i}(V_{n-1}),\quad n\geq 1,$
with $V_{*}=\bigcup_{n\geq 1}V_{n}$. Following (1.2), denote the vertices $V_{n}=\{v_{1},v_{2},v_{3},\dots\}$. Define the Dirichlet energies
$\mathcal{E}_{n}(u)=\sum_{i,j=1}^{|V_{n}|}c_{ij}^{n}\frac{(u(v_{j})-u(v_{i}))^{2}}{2},\quad u\in L(V_{n},\mathbb{R}),$ (7.2)
where $c_{ij}^{n}\geq 0$. As before, it is possible to extend the domain of $\mathcal{E}_{n}$ to all $u\in C(K,\mathbb{R})$. We assume that the $\mathcal{E}_{n}$ satisfy the following two conditions.
  1. 1.
    There are $0<r_{i}<1$ so that
    $\mathcal{E}_{n}(u)=\sum_{i=1}^{N}r_{i}^{-1}\mathcal{E}_{n-1}(u\circ F_{i}),\quad u\in L(V_{n},\mathbb{R}).$ (7.3)
  2. 2.
    For every $n\geq 1$,
    $\mathcal{E}_{n-1}(u)=\min\{\mathcal{E}_{n}(\bar{u})\colon\bar{u}\in L(V_{n},\mathbb{R}),\bar{u}|_{V_{n-1}}=u\}.$ (7.4)
In particular, (7.4) implies that $\mathcal{E}_{n}(u)$ is non-decreasing for any $u\in C(K,\mathbb{R})$ so the limit
$\mathcal{E}(u)=\lim_{n\to\infty}\mathcal{E}_{n}(u)$
is well-defined. Moreover, it is possible to define a harmonic extension algorithm analogous to (2.9) [18]. As before write
$\operatorname{dom}(\mathcal{E})=\{u\in C(K,\mathbb{R})\colon\mathcal{E}(u)<\infty\}.$
The energies (7.2) also naturally induce graph structures $\Gamma_{n}=(V_{n},E_{n})$. Namely, for all $v_{j},v_{i}\in V_{n}$,
$i\sim_{n}j\mbox{ iff }c_{ij}^{n}>0.$
To define the degree of a $\mathbb{T}$-valued map on a p.c.f. fractal $K$, we must describe the cycle space, $Cyc(\Gamma_{m})$ of the approximating graph $\Gamma_{m}$ of $K$. For $SG$, the boundaries of the $n$-cells $\partial F_{w}(K)$ sufficed. In general, however, there is no canonical way to form a basis for this cycle space, so one must choose it. Let
$\mathcal{B}_{m}=\left\{\gamma_{1}^{(m)},\dots,\gamma_{n_{m}}^{(m)}\right\}$ (7.5)
be a basis for $Cyc(\Gamma_{m})$. For $f\in C(K,\mathbb{T})$, define the degree
$\bar{\omega}^{(m)}(f)=\left(\omega_{\gamma_{1}^{(m)}}(f),\dots,\omega_{\gamma_{n_{m}}^{(m)}}(f)\right),$
where $\omega_{\gamma}(f)$ is the degree of $f|_{\gamma}$.
Again, to generate harmonic maps, we construct a covering space depending on the degree. To that end, we require a few technical assumptions, described below. We believe they are not overly restrictive, e.g., they hold for standard examples of p.c.f. fractals such as the higher order Sierpinski Gaskets, the pentagasket, and the hexagasket [14].
Fix an embedding $\iota:\mathcal{B}_{m}\to\operatorname{Cyc}(\Gamma_{m+1})$, such that
$V\left(\gamma^{(m)}_{i}\right)\subset V\left(\iota(\gamma^{(m)}_{i})\right),\quad 1\leq i\leq n_{m},$ (7.6)
and
$V\left(\iota(\gamma^{(m)}_{i})\right)\cap V\left(\iota(\gamma^{(m)}_{j})\right)=V\left(\gamma^{(m)}_{i}\right)\cap V\left(\gamma^{(m)}_{j}\right),\quad 1\leq i,j\leq n_{m},\;\;\;{j\neq i}.$ (7.7)
Assume that for every $i\in[n_{m}]$ there is $\xi(\gamma_{i}^{(m)})\in V_{m+1}$ such that
$V\left(\iota(\gamma^{(m)}_{i})\right)\ni\xi(\gamma_{i}^{(m)})\notin V\left(\iota(\gamma^{(m)}_{j})\right),\quad j\neq i.$ (7.8)
In particular, the $\xi(\gamma_{i}^{(m)})$ serve the role of the cut vertices.
From (7.6)-(7.8), we can construct a covering space $\tilde{K}$ using the same procedure as in Section 5. Indeed, first fix a degree
$\bar{\omega}^{(m)}(f)=(\rho_{1},\dots,\rho_{n_{m}})\in\mathbb{Z}^{n_{m}}.$ (7.9)
Let $K_{\times}=K\times\mathbb{Z}$. On each sheet $K^{s}=K\times\{s\}$, create two copies of the cut vertices $z_{i}^{s}:=\xi(\gamma_{i}^{(m)})$, named $(z_{i}^{s})_{-}$ and $(z_{i}^{s})_{+}$. Identifying
$(z_{i}^{s})_{+}\simeq(z_{i}^{s+\rho_{i}})_{-},$
results in the covering space $\tilde{K}=K_{\times}/\simeq$ with fundamental domain $K^{0}$. Additionally, $K^{0}$ is approximated by graphs $\Gamma_{m}^{s}=(V_{m}^{s},E_{n}^{s})$.
As a visual example, Figure 12 shows this construction on $\Gamma_{1}$ of $SG_{3}$, a higher-order Sierpinski Gasket (see, e.g., [14] for a precise construction of this higher order SG). Given degree $\bar{\omega}^{(m)}$, gluing cut points across sheets results in the associated covering space $\tilde{K}$.
a)  Refer to caption      b)  Refer to caption
Figure 12: Covering space construction for the higher-order Sierpinski Gasket, $SG_{3}$. a) A basis for the cycle space of $\Gamma_{1}$. In particular, the boundaries of the $n$-cells are not sufficient to form a basis. b) Associated cut vertices on $\Gamma_{2}$. Superscripts omitted for simplicity.
7.2 Convergence of KM on p.c.f. fractals
An analogous theory of harmonic functions exists on these covering spaces. Introduce Dirichlet energies on the fundamental domain,
${\mathcal{E}}_{n}[\Gamma_{n}^{0}](u)=\sum_{(i,j)\in E^{0}_{n}}c_{ij}^{n}\frac{(u(v_{j})-u(v_{i}))^{2}}{2},$
and corresponding spaces
$H^{0}_{n}=\{f\in L(V^{0}_{n},\mathbb{R})\colon$ $f(v^{0}_{1})=0,$
$\;f((z^{0}_{i})_{+})=f((z^{0}_{i})_{-})+\rho_{i},\forall i\in[n_{m}]\},$
where, as usual, the value at $v_{1}^{0}\in V^{0}$ is fixed to remove translation invariance. As in (5.8), the limit
${\mathcal{E}}_{K^{0}}(f)=\lim_{n\to\infty}{\mathcal{E}}_{n}[\Gamma_{n}^{0}](f)$
is well-defined. Denote $\operatorname{dom}({\mathcal{E}}_{K^{0}})\doteq\{u\in C(K^{0},\mathbb{R}):{\mathcal{E}}_{K^{0}}(u)<\infty\}$.
Minimizing ${\mathcal{E}}_{n}[\Gamma_{n}^{0}]$ and repeatedly applying the harmonic extension algorithm results in a harmonic function on $K^{0}$, which can be extended to all of $\tilde{K}$ as in Section 5.2. Projecting the range to $\mathbb{T}$ results in the desired harmonic map. Moreover, the resultant map satisfies natural boundary conditions (cf. Lemma 5.3).
Theorem 7.1.
Fix $m\in\mathbb{N}$ and a degree
$\bar{\omega}^{(m)}(f)=(\rho_{1},\dots,\rho_{n_{m}}).$ (7.10)
There is a harmonic map $f\colon K\to\mathbb{T}$ satisfying (7.10) and $\partial_{\bf n}f(v_{i})=0$ for all $v_{i}\in V^{0}$.
As in the previous sections, to establish convergence results we require Hölder continuity. A variation of the following can be found in [18]. Recall that the fundamental domain $K^{0}$ can be represented as a finite union of $m$-cells. Denoting $K^{0}_{w}\doteq K_{w}(K^{0})$, we have $K^{0}=\bigcup_{|w|=m}K^{0}_{w}$, where each $K^{0}_{w}$ is a copy of $K$.
Lemma 7.2.
Let $f\in\operatorname{dom}(\mathcal{E}_{K^{0}})$, and represent the fundamental domain as a union of $m$-cells: $K^{0}=\bigcup_{|w|=m}K^{0}_{w}$. There exists $\beta>0$ so that for any $x,y\in K^{0}_{w}$,
$\frac{|f(x)-f(y)|}{|x-y|^{\beta}}\leq C\sqrt{\mathcal{E}_{K^{0}}(f)},$
where $C$ is independent of $f$.
Proof.
Arguing as in Lemma 5.4, it is sufficient to establish Hölder continuity on $K$ itself. Let $R(x,y)$ be the effective resistance metric induced by ${\mathcal{E}}_{K}$ on $K$. That is,
$R(x,y)^{-1}\doteq\min\{{\mathcal{E}}(u)\colon u(x)=0,\;u(y)=1\}.$
Then $R$ is a metric and induces a 1/2-Hölder norm on the domain of ${\mathcal{E}}_{K}$, cf [10]:
$|f(x)-f(y)|\leq R(x,y)^{1/2}\sqrt{{\mathcal{E}}(f)}.$
Hence, we need only show that $R(x,y)^{1/2}\leq C|x-y|^{\beta}$.
To that end, assume first that $x,y\in V_{n}$ and take $u\in\operatorname{dom}({\mathcal{E}})$ to be the harmonic spline whose values on $V_{n}$ are $u(z)=\delta_{zy}$ (i.e., $u(z)=1$ only when $z=y$). Then ${\mathcal{E}}(u)\geq{\mathcal{E}}_{n}(u)\geq Cr_{w}^{-1}$, where $r_{w}$ is a product of the $r_{i}$ defined in (7.3). From the definition of $R$ certainly $R(x,y)\leq Cr_{w}$. Defining $\bar{r}=\max r_{i}<1$, we have $R(x,y)\leq C\bar{r}^{n}$. Next, recall that the $F_{i}$ have contraction ratios $\lambda_{i}$. Letting $\underline{\lambda}=\min\lambda_{i}$, we have a lower bound $|x-y|\geq C\underline{\lambda}^{n}$, $x,y\in V_{n}$. Thus taking $\alpha=\log(\overline{r})/\log(\underline{\lambda})$ then
$R(x,y)^{1/2}\leq C|x-y|^{\alpha/2}.$
If $x,y\in K$ are arbitrary, then by density of $V_{*}$ there are sequences $x_{n}\to x$ and $y_{n}\to y$ with $x_{n},y_{n}\in V_{n}$ and the same estimate follows by continuity. ∎
The convergence of stable equilibria of KM on p.c.f. fractals now follows similarly. Consider the KM on $\Gamma_{n}$ approximating $K$:
$\dot{u}(t,v_{i})=\sum_{i,j}c_{i,j}^{n}\sin\left(2\pi\left(u(t,v_{j})-u(t,v_{i})\right)\right),\quad v_{i}\in V_{n}.$ (7.11)
Combining the covering space construction of harmonic maps (Theorem 7.1), and the a priori Hölder estimates on the fundamental domain (Lemma 7.2), applying the same arguments as in Section 6 results in the following Theorem.
Theorem 7.3.
Fix $m\in\mathbb{N}$ and a degree $\bar{\omega}^{(m)}(u^{*})$. Construct a harmonic map $u^{*}\in C(K,\mathbb{T})$ as in Theorem 7.1. Then, for each $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for all $n\geq N$, the KM on $\Gamma_{n}$ has a stable steady state solution $u^{n}\in L(V_{n},\mathbb{T})$ such that
$\max_{x\in V_{n}}|u^{n}(x)-u^{*}(x)|<\epsilon.$ (7.12)
Moreover, there is an extension of $u^{n}$ to a continuous function on $K$, $\tilde{u}^{n}$ such that $\bar{\omega}^{(m)}(\tilde{u}^{n})=\bar{\omega}^{(m)}(u^{*})$ and
$\max_{x\in K}|\tilde{u}^{n}(x)-u^{*}(x)|<\epsilon.$ (7.13)
Author Contributions  G.S.M. and M.S.M. contributed to writing the manuscript equally. All authors reviewed the manuscript.
Funding  The work of G.S.M and M.S.M. was partially supported by NSF DMS Awards # 2406941 and
# 2406942 respectively.
Data availability  No datasets were generated or analysed during the current study.